# Markovian Promoter Models: A Mechanistic Alternative to Hill Functions in Gene Regulatory Networks

### v3

@@ 2.7 Detailed Mathematical Conversion: GAL System
We present a comprehensive, step-by-step conversion of the yeast galactose (GAL) regulatory system from the traditional ODE formulation with Hill functions [ramsey2006dual] to our Markovian promoter framework. This detailed exposition serves as a template for converting other gene regulatory systems.
@@ 2.7.1 Step 1: Original ODE Model
The GAL system regulates galactose metabolism in Saccharomyces cerevisiae through a network of five genes: GAL1, GAL2, GAL3, GAL80, and GAL4 [lohr1995transcriptional, plattYeastGalactoseGenetic1998]. The original ODE model from Ramsey et al. (2006) [ramsey2006dual] uses Hill functions to describe transcriptional regulation.
@@
Similarly, for GAL2 (5 binding sites), GAL3 (1 binding site), and GAL80 (1 binding site), the functions follow standard Hill kinetics as detailed in the original model [ramsey2006dual].
@@ 2.7.10 Step 10: Validation and Comparison
The Markovian model results are validated by comparing with the original ODE model for dose-response behavior, the full SSA for stochastic trajectories and noise characteristics, and experimental data from literature [ramsey2006dual, venturelli2012synergistic]. We expect quantitative agreement in mean trajectories and noise levels, with the Markovian model providing a significant computational speedup over SSA.
@@ 2.8.1 Repressilator: Cyclic Repression Network
The repressilator [elowitz2000synthetic] consists of three genes (lacI, tetR, cI) in a cyclic repression topology where each gene represses the next.
@@ 2.8.2 Goodwin Oscillator: Negative Feedback with High Cooperativity
The Goodwin oscillator [goodwin1965oscillatory] is a minimal model for circadian rhythms characterized by strong negative feedback.
@@ 2.8.3 Toggle Switch: Mutual Repression and Bistability
The toggle switch [gardner2000construction] consists of two genes that mutually repress each other, creating bistability.
@@ 2.8.4 Incoherent Feed-Forward Loop (I1-FFL): Multi-Input Logic
The I1-FFL [mangan2003structure] features a master regulator X that activates both an intermediate Y and a target Z, while Y represses Z.
@@
The p53-Mdm2 system [lahav2004dynamics] is a negative feedback loop where p53 activates Mdm2 transcription, and Mdm2 promotes p53 degradation.
@@ 2.9.3 Convergence Analysis
2.9.5 Adaptive Time Stepping (Future Work)
@@ 2.9.7 Implementation Notes
3 Results
@@ 3.2 GAL System: Comprehensive Four-Way Comparison
The yeast galactose (GAL) regulatory system serves as our primary validation case due to its well-characterized kinetics [ramsey2006dual, venturelli2012synergistic] and variable promoter architecture (1-5 binding sites across different genes).
@@ 3.2.1 Dose-Response Curves
The dose-response curves demonstrate several key features. First, ultrasensitivity emerges naturally from discrete states. Despite using simple mass-action kinetics for individual binding events, the Markovian model reproduces the ultrasensitive dose-response curves characteristic of cooperative binding. For GAL1 (4 binding sites), the Hill coefficient estimated from the Markovian model ($n_{eff}=3.2\pm 0.3$) closely matches both the ODE model ($n=4$) and experimental measurements [ramsey2006dual]. This demonstrates that cooperativity emerges from the combinatorial state space without requiring phenomenological Hill functions.
@@ 3.2.2 Time Series Dynamics
Furthermore, stochastic fluctuations match SSA predictions. We quantified the coefficient of variation (CV) for each gene at steady state (Table 6). The Markovian model reproduces SSA noise levels with high fidelity, with CV ratios of 0.95-0.96. This demonstrates that promoter-level stochasticity is the dominant source of gene expression noise in this system [raj2008nature, sanchez2013regulation].
@@ 3.2.3 Promoter State Dynamics
Figure 3: Promoter state dynamics for GAL1 gene. (A) Heatmap showing probability distribution over all 15 promoter states (rows) as a function of time (columns). State (k,m) indicates k Gal4p bound and m Gal80p bound. Color intensity represents probability. (B) Single stochastically sampled promoter state trajectory showing discrete jumps between states. (C) Gal4p dimer (blue) and Gal80p dimer (orange) concentrations driving promoter transitions. Following galactose induction at t=10 min, the promoter rapidly transitions from empty state (0,0) to active states (k¿m), then partially to repressed states (k=m) as Gal80p accumulates. The equilibrium distribution at t¿100 min shows predominance of partially active states (2-3 Gal4p bound, 0-1 Gal80p bound), consistent with experimental ChIP data [venturelli2012synergistic].
@@ 3.2.4 Comparison with Experimental Data
We compared model predictions with experimental measurements from Ramsey et al. [ramsey2006dual] for protein fold-changes upon galactose induction (Table 7).
@@ 3.3.1 Repressilator: Sustained Oscillations from Cyclic Repression
The repressilator [elowitz2000synthetic] is a landmark synthetic biology circuit consisting of three genes in a ring of mutual repression. It serves as a test case for oscillatory dynamics and stochastic phase variations.
@@ 3.3.2 Goodwin Oscillator: High Cooperativity and Negative Feedback
The Goodwin oscillator [goodwin1965oscillatory] represents circadian rhythm dynamics with strong negative feedback ($n\sim 10$).
@@ 3.3.3 Toggle Switch: Bistability and Rare Switching
The toggle switch [gardner2000construction] exhibits bistability through mutual repression, with rare stochastic transitions between stable states.
@@ 3.4.1 Incoherent Feed-Forward Loop: Multi-Input Logic and Pulse Generation
The I1-FFL is a ubiquitous network motif [mangan2003structure, alon2007network] that implements temporal filtering through combinatorial logic. It tests the multi-input promoter framework.
@@
The I1-FFL pulse generation has been experimentally characterized in the E. coli arabinose system [mangan2006structure]. Our model predictions agree with experimental observations. The model predicts a pulse duration of 30 min compared to 25-35 min in experiments, a fold-change of 8-10× compared to 7-12×, and peak timing at roughly 25 min compared to 20-30 min.
@@
The p53-Mdm2 system [lahav2004dynamics, geva2006dynamics] exhibits damped oscillations in response to DNA damage, with p53 activating Mdm2 transcription and Mdm2 promoting p53 degradation.
@@
The Mdm2 promoter has 2 p53 binding sites with cooperative binding ($q_{r}=5$). We analyzed the promoter state distribution at different p53 levels. Cooperative binding manifests as distinct dominant states at different concentrations: state 0 (no binding) dominates at low p53 (<< 2 nM), mixed states appear at intermediate p53 (2-5 nM), and state 2 (both sites bound) dominates at high p53 (>> 5 nM). The effective Hill coefficient from the Markovian model ($n_{eff}=1.8\pm 0.2$) matches experimental measurements ($n_{exp}=1.7-2.0$) [wu1993mdm2], validating the cooperative binding implementation.
@@
Lahav et al. [lahav2004dynamics] measured p53 oscillations in single cells using fluorescent reporters. Our Markovian model predictions agree with their observations. We predict a period of 5.5 h compared to 5.2 ± 0.9 h in experiments, cell-to-cell variability (CV) of 0.14 compared to 0.12-0.16, and an average of 3.7 pulses compared to 3-4 observed experimentally.
@@ 3.6.3 Sources of Gene Expression Noise
We decomposed the total noise into contributions from different sources using the noise decomposition framework [paulsson2004summing, sanchez2013regulation]. For the Repressilator, promoter switching contributes 49% of the noise, while translation contributes 31%, with transcription and degradation combining for the remaining 20% (Table 17). This reveals that capturing only promoter noise while ignoring translation noise results in 68% underestimation of total noise, explaining the low CV observed in the promoter-only stochastic model.
@@ 3.6.7 Summary: Protein Stochasticity
4 Discussion
4.1 Enabling Gene Regulation in Whole-Cell Models
@@ 4.1.2 Preserving Stochasticity
Unlike deterministic Hill functions, Markovian promoters capture three distinct sources of gene expression noise. First, promoter switching noise arises from the stochastic transitions between active and inactive promoter states, creating transcriptional bursting. Second, transcriptional noise reflects the Poisson statistics of mRNA production events given a promoter state. Third, translational noise accounts for the variability in protein production events. Capturing these noise sources is essential for whole-cell models where stochastic effects can dominate cellular decision-making, particularly for low-copy-number genes [elowitz2002stochastic].
@@ 4.1.3 Practical Parameterization
A major advantage for whole-cell modeling is that parameters can be inferred from accessible experimental data, circumventing the need for difficult-to-obtain kinetic rates. Transcriptomic RNA-seq data provides steady-state mRNA levels that enable the inference of effective dissociation constants and transcription rates. Proteomics data supplies protein levels and degradation rates. Furthermore, noise characteristics derived from single-cell RNA-seq can constrain promoter switching rates, while ChIP-seq data validates binding site numbers. This contrasts with detailed kinetic modeling, where individual $k_{on}$ and $k_{off}$ values are rarely known, typically requiring extensive in vitro binding assays. Our coarse-grained approach bridges this gap by mapping binding kinetics to observable omics data.
@@ 4.5.2 Computational Scaling
4.5.3 Example: E. coli Whole-Cell Model
Consider integrating gene regulation into an E. coli whole-cell model with approximately 200 transcription factors and 4,000 genes, of which roughly 1,000 have known regulation. A full explicit binding model would require defining unique species for every possible TF-DNA complex, potentially adding around 200,000 species and 400,000 reactions, rendering the simulation computationally intractable. In contrast, the Markovian promoter model adds only about 1,000 promoter state variables and 4,000 reactions (assuming an average of two binding sites per regulated promoter). This level of complexity is well within the feasibility range of current simulation hardware.

### v4

@@ 0.7 Detailed Mathematical Conversion: GAL System
We present a comprehensive, step-by-step conversion of the yeast galactose (GAL) regulatory system from the traditional ODE formulation with Hill functions (Ramsey et al., 2006) to our Markovian promoter framework. This detailed exposition serves as a template for converting other gene regulatory systems.
@@ 0.7.1 Step 1: Original ODE Model
The GAL system regulates galactose metabolism in Saccharomyces cerevisiae through a network of five genes: GAL1, GAL2, GAL3, GAL80, and GAL4 (Lohr et al., 1995; Platt and Reece, 1998). The original ODE model from Ramsey et al. (2006) (Ramsey et al., 2006) uses Hill functions to describe transcriptional regulation.
@@
Similarly, for GAL2 (5 binding sites), GAL3 (1 binding site), and GAL80 (1 binding site), the functions follow standard Hill kinetics as detailed in the original model (Ramsey et al., 2006).
@@ 0.7.10 Step 10: Validation and Comparison
The Markovian model results are validated by comparing with the original ODE model for dose-response behavior, the full SSA for stochastic trajectories and noise characteristics, and experimental data from literature (Ramsey et al., 2006; Venturelli et al., 2012). We expect quantitative agreement in mean trajectories and noise levels, with the Markovian model providing a significant computational speedup over SSA.
@@ 0.8.1 Repressilator: Cyclic Repression Network
The repressilator (Elowitz and Leibler, 2000) consists of three genes (lacI, tetR, cI) in a cyclic repression topology where each gene represses the next.
@@ 0.8.2 Goodwin Oscillator: Negative Feedback with High Cooperativity
The Goodwin oscillator (Goodwin, 1965) is a minimal model for circadian rhythms characterized by strong negative feedback.
@@ 0.8.3 Toggle Switch: Mutual Repression and Bistability
The toggle switch (Gardner et al., 2000) consists of two genes that mutually repress each other, creating bistability.
@@ 0.8.4 Incoherent Feed-Forward Loop (I1-FFL): Multi-Input Logic
The I1-FFL (Mangan and Alon, 2003) features a master regulator X that activates both an intermediate Y and a target Z, while Y represses Z.
@@
The p53-Mdm2 system (Lahav et al., 2004) is a negative feedback loop where p53 activates Mdm2 transcription, and Mdm2 promotes p53 degradation.
@@ 0.9.3 Convergence Analysis
0.9.5 Adaptive Time Stepping
@@ 0.9.7 Implementation Notes
0.10 ChEC-seq Markovian Dwell Time Model
To mechanistically bridge static genomic binding data with dynamic transcription kinetics, we developed a Sequential Transcription Assembly Model that infers kinetic rates directly from ChEC-seq dwell time measurements.
0.10.1 ChEC-seq Dwell Time Derivation
ChEC-seq (Chromatin Endogenous Cleavage with high-throughput sequencing) provides a quantitative measure of protein residency at specific genomic loci (Rossi et al., 2021). Unlike traditional ChIP-seq, the time-resolved nature of ChEC-seq cleavage allows us to interpret signal intensity as a proxy for the fractional residence time (dwell time) of a factor at the promoter (VanBelzen et al., 2024). For a given gene $g$, we normalize the integrated signal intensities $I_{p}$ for each protein $p$ in the assembly complex (TF, Cofactor, TFIID, RNAPII) to obtain fractional dwell times $\tau_{p}$:
$\tau_{p}=\frac{I_{p}}{\sum_{j\in\text{complex}}I_{j}}$ (5)
This normalization assumes that the total signal represents the complete assembly cycle, and $\tau_{p}$ reflects the proportion of the cycle during which protein $p$ is chemically cross-linked and present.
0.10.2 Sequential Markovian Formulation
We model the promoter transition as a linear chain of $N$ discrete states, representing the ordered recruitment of the transcription machinery:
  1. 1.
    Empty Promoter
  2. 2.
    TF Bound
  3. 3.
    Cofactor Recruitment
  4. 4.
    TFIID/GTF Recruitment
  5. 5.
    RNAPII Recruitment (Pre-Initiation Complex)
  6. 6.
    Elongation
Transitions between state $i$ and $i+1$ are governed by forward rate $k_{on,i}\cdot[P_{i}]$ and backward rate $k_{off,i}$. The final elongation step is modeled as an irreversible transition that resets the cycle. The steady-state probability distribution $\mathbf{\pi}$ is computed by solving $\mathbf{\pi}\mathbf{Q}=0$, where $\mathbf{Q}$ is the rate matrix.
0.10.3 Optimization Procedure
We estimate the kinetic parameters $\theta=\{k_{on},k_{off}\}$ for each gene by minimizing the discrepancy between the model-predicted fractional residence times and the experimentally derived dwell times $\tau$. The objective function is:
$J(\theta)=\sum_{p}\left(\hat{\tau}_{p}(\theta)-\tau_{p,\text{exp}}\right)^{2}+\lambda\mathcal{P}(\theta)$ (6)
where $\hat{\tau}_{p}(\theta)$ is the sum of steady-state probabilities for all states containing protein $p$, normalized by the total occupation. The penalty term $\mathcal{P}(\theta)$ enforces biological constraints, ensuring that:
  • Forward association rates favor complex formation ($k_{on}>k_{off}/100$).
  • Recruitment rates decrease realistically for larger complexes.
This optimization yields a gene-specific kinetic model that is quantitatively consistent with the observed residence times of the transcriptional machinery.
Results
@@ 0.12 GAL System: Comprehensive Four-Way Comparison
The yeast galactose (GAL) regulatory system serves as our primary validation case due to its well-characterized kinetics (Ramsey et al., 2006; Venturelli et al., 2012) and variable promoter architecture (1-5 binding sites across different genes).
@@ 0.12.1 Dose-Response Curves
The dose-response curves demonstrate several key features. First, ultrasensitivity emerges naturally from discrete states. Despite using simple mass-action kinetics for individual binding events, the Markovian model reproduces the ultrasensitive dose-response curves characteristic of cooperative binding. For GAL1 (4 binding sites), the Hill coefficient estimated from the Markovian model ($n_{eff}=3.2\pm 0.3$) closely matches both the ODE model ($n=4$) and experimental measurements (Ramsey et al., 2006). This demonstrates that cooperativity emerges from the combinatorial state space without requiring phenomenological Hill functions.
@@ 0.12.2 Time Series Dynamics
Furthermore, stochastic fluctuations match SSA predictions. We quantified the coefficient of variation (CV) for each gene at steady state (Table 6). The Markovian model reproduces SSA noise levels with high fidelity, with CV ratios of 0.95-0.96. This demonstrates that promoter-level stochasticity is the dominant source of gene expression noise in this system (Raj and Van Oudenaarden, 2008; Sanchez and Golding, 2013).
@@ 0.12.3 Promoter State Dynamics
Figure 3: Promoter state dynamics for GAL1 gene. (A) Heatmap showing probability distribution over all 15 promoter states (rows) as a function of time (columns). State (k,m) indicates k Gal4p bound and m Gal80p bound. Color intensity represents probability. (B) Single stochastically sampled promoter state trajectory showing discrete jumps between states. (C) Gal4p dimer (blue) and Gal80p dimer (orange) concentrations driving promoter transitions. Following galactose induction at t=10 min, the promoter rapidly transitions from empty state (0,0) to active states (k¿m), then partially to repressed states (k=m) as Gal80p accumulates. The equilibrium distribution at t¿100 min shows predominance of partially active states (2-3 Gal4p bound, 0-1 Gal80p bound), consistent with experimental ChIP data (Venturelli et al., 2012).
@@ 0.12.4 Comparison with Experimental Data
We compared model predictions with experimental measurements from Ramsey et al. (Ramsey et al., 2006) for protein fold-changes upon galactose induction (Table 7).
@@ 0.13.1 Repressilator: Sustained Oscillations from Cyclic Repression
The repressilator (Elowitz and Leibler, 2000) is a landmark synthetic biology circuit consisting of three genes in a ring of mutual repression. It serves as a test case for oscillatory dynamics and stochastic phase variations.
@@ 0.13.2 Goodwin Oscillator: High Cooperativity and Negative Feedback
The Goodwin oscillator (Goodwin, 1965) represents circadian rhythm dynamics with strong negative feedback ($n\sim 10$).
@@ 0.13.3 Toggle Switch: Bistability and Rare Switching
The toggle switch (Gardner et al., 2000) exhibits bistability through mutual repression, with rare stochastic transitions between stable states.
@@ 0.14.1 Incoherent Feed-Forward Loop: Multi-Input Logic and Pulse Generation
The I1-FFL is a ubiquitous network motif (Mangan and Alon, 2003; Alon, 2007) that implements temporal filtering through combinatorial logic. It tests the multi-input promoter framework.
@@
The I1-FFL pulse generation has been experimentally characterized in the E. coli arabinose system (Mangan et al., 2006). Our model predictions agree with experimental observations. The model predicts a pulse duration of 30 min compared to 25-35 min in experiments, a fold-change of 8-10× compared to 7-12×, and peak timing at roughly 25 min compared to 20-30 min.
@@
The p53-Mdm2 system (Lahav et al., 2004; Geva-Zatorsky et al., 2006) exhibits damped oscillations in response to DNA damage, with p53 activating Mdm2 transcription and Mdm2 promoting p53 degradation.
@@
The Mdm2 promoter has 2 p53 binding sites with cooperative binding ($q_{r}=5$). We analyzed the promoter state distribution at different p53 levels. Cooperative binding manifests as distinct dominant states at different concentrations: state 0 (no binding) dominates at low p53 (<< 2 nM), mixed states appear at intermediate p53 (2-5 nM), and state 2 (both sites bound) dominates at high p53 (>> 5 nM). The effective Hill coefficient from the Markovian model ($n_{eff}=1.8\pm 0.2$) matches experimental measurements ($n_{exp}=1.7-2.0$) (Wu et al., 1993), validating the cooperative binding implementation.
@@
Lahav et al. (Lahav et al., 2004) measured p53 oscillations in single cells using fluorescent reporters. Our Markovian model predictions agree with their observations. We predict a period of 5.5 h compared to 5.2 ± 0.9 h in experiments, cell-to-cell variability (CV) of 0.14 compared to 0.12-0.16, and an average of 3.7 pulses compared to 3-4 observed experimentally.
@@ 0.16.3 Sources of Gene Expression Noise
We decomposed the total noise into contributions from different sources using the noise decomposition framework (Paulsson, 2004; Sanchez and Golding, 2013). For the Repressilator, promoter switching contributes 49% of the noise, while translation contributes 31%, with transcription and degradation combining for the remaining 20% (Table 17). This reveals that capturing only promoter noise while ignoring translation noise results in 68% underestimation of total noise, explaining the low CV observed in the promoter-only stochastic model.
@@ 0.16.7 Summary: Protein Stochasticity
0.17 Validation of ChEC-seq Markovian Dwell Time Model
0.17.1 Reproduction of Assembly Dynamics
We applied the Sequential Assembly Dwell Time Model to a set of representative yeast genes. By fitting the kinetic parameters to ChEC-seq data, the model successfully reproduces the observed fractional residence times of key assembly factors (TF, Mediator, TFIID, RNAPII).
Figure 11: Validation of Sequential Assembly Dwell Time Model. Correlation between experimentally derived fractional residence times (ChEC-seq) and model-predicted occupancies for key transcriptional components (Pearson $r>0.85$). Each point represents a protein factor for a specific gene.
Figure 11 shows the strong correlation (Pearson $r>0.85$) between the experimental ChEC-seq dwell times and the model-predicted residency fractions across the test set. This agreement confirms that a linear sequential Markovian framework is sufficient to capture the steady-state occupancy profiles revealed by chromatin profiling.
0.17.2 Insights into Complex Stability
The optimized kinetic parameters provide mechanistic insights into the stability of assembly intermediates, revealing distinct regulatory strategies. We consistently observed that the transition from Cofactor recruitment to TFIID recruitment represents a significant kinetic bottleneck, characterized by high $k_{off}$ relative to $k_{on}$ in genes with low basal expression. In contrast, highly expressed genes displayed ”trap-like” kinetics where downstream recruitment steps (TFIID $\to$ RNAPII) became increasingly irreversible, effectively funneling the system toward productive elongation (VanBelzen et al., 2024). This analysis suggests that transcriptional bursting patterns are essentially encoded in the stability of these intermediate complexes, a feature that is quantitatively extractable from static dwell time measurements using our modeling framework. These findings align with recent single-molecule studies suggesting that PIC stability is a primary determinant of burst frequency (Larson et al., 2011).
Discussion
@@ 0.17.4 Preserving Stochasticity
Unlike deterministic Hill functions, Markovian promoters capture three distinct sources of gene expression noise. First, promoter switching noise arises from the stochastic transitions between active and inactive promoter states, creating transcriptional bursting. Second, transcriptional noise reflects the Poisson statistics of mRNA production events given a promoter state. Third, translational noise accounts for the variability in protein production events. Capturing these noise sources is essential for whole-cell models where stochastic effects can dominate cellular decision-making, particularly for low-copy-number genes (Elowitz et al., 2002).
@@ 0.17.5 Practical Parameterization
A major advantage for whole-cell modeling is that parameters can be inferred from accessible experimental data, circumventing the need for difficult-to-obtain kinetic rates. Transcriptomic RNA-seq data provides steady-state mRNA levels that enable the inference of effective dissociation constants and transcription rates. Proteomics data supplies protein levels and degradation rates. Furthermore, noise characteristics derived from single-cell RNA-seq can constrain promoter switching rates, while ChIP-seq data validates binding site numbers.
Meanwhile, we introduce a method to parameterize short-lived assembly intermediates using ChEC-seq dwell times. By equating the normalized ChEC-seq signal to the steady-state probability of Markovian states, we can mathematically invert the master equation to solve for the microscopic rate constants ($k_{on},k_{off}$) that generate the observed occupancy profiles. This effectively allows us to ”read” dynamic assembly kinetics from static genome-wide binding data.
This contrasts with detailed kinetic modeling, where individual $k_{on}$ and $k_{off}$ values are rarely known, typically requiring extensive in vitro binding assays. Our coarse-grained approach bridges this gap by mapping binding kinetics to observable omics data.
@@ 0.17.6 Handling Nonlinearity Without Breaking Mass-Action
0.18 Enabling Gene Regulation in Whole-Cell Models
@@ 0.22.2 Computational Scaling
0.22.3 Example: Yeast Whole-Cell Model
Consider integrating transcriptional regulation into a budding yeast (Saccharomyces cerevisiae S288C) whole-cell model. The yeast genome contains approximately 6,604 ORFs bionumbers_yeast_genes_6604, and encodes on the order of $\sim$200 sequence-specific transcription factors. If regulation were implemented using a fully explicit TF–DNA binding formulation, one would need to introduce a distinct molecular species for every TF–promoter complex. Even if only $\sim 1{,}000$ promoters have curated TF regulation, this would still require on the order of $O(N_{\mathrm{TF}}\times N_{\mathrm{promoter}})\approx 2\times 10^{5}$ TF–DNA complex species and $\gtrsim 4\times 10^{5}$ binding/unbinding reactions, quickly rendering stochastic whole-cell simulations computationally intractable.
In contrast, a Markovian promoter-state approximation scales linearly with the number of regulated promoters. For example, modeling $\sim$1,000 regulated promoters with an average of two effective binding sites per promoter adds only $\sim$1,000 promoter state variables and $\sim$4,000 transition reactions, remaining well within the feasible range of modern simulation hardware.